The equation B=1+3n tells the number of beams required to build a frame of n sections. How many sections can be built with 79 beams?
Answer by tutorcecilia(2152)

B=1+3n
79=1+3n [substitute 79 for the B-variable]
79-1=1+3n [solve for the n-variable]
78=3n
78/3=3n/3
26 sections=n
Checking:
79=1+3(26)
79=79